# Distillation of Discrete Diffusion through Dimensional Correlations

### v3

@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b) have demonstrated excellent performance in generative modeling, particularly for continuous data such as images (Nichol et al., 2021; Rombach et al., 2022; Saharia et al., 2022), audio (Kong et al., 2020; Chen et al., 2021; Evans et al., 2024), and video (Harvey et al., 2022; Ho et al., 2022; Blattmann et al., 2023). Recent advancements in diffusion models often outperform traditional generative models, such as variational autoencoders (VAEs, Kingma & Welling, 2014; Higgins et al., 2017; Zhao et al., 2019) and generative adversarial networks (GANs, Goodfellow et al., 2014), in terms of sample quality and the controllability of the generated results. Furthermore, diffusion models are not limited to learning continuous data; they can also be applied to discrete or categorical data with modifications (Hoogeboom et al., 2021; Austin et al., 2021) and offer a promising approach for discrete generative modeling (Gu et al., 2022; Lou et al., 2024). Such discrete diffusion models are the main topic of this paper.
A notable drawback of diffusion models is their slow sampling speed due to requiring many sampling steps (Xiao et al., 2022; Zhang & Chen, 2023). In continuous domains, various approaches have been proposed to reduce the number of steps, including well-designed forward processes (Song et al., 2021a) and fast solvers of stochastic/ordinary differential equations (SDEs/ODEs) (Lu et al., 2022a, b; Zheng et al., 2023b). Another notable approach is knowledge distillation, which significantly reduces the number of sampling steps compared with earlier attempts by compressing pretrained diffusion models into single- or few-step generative models (Luhman & Luhman, 2021; Salimans & Ho, 2022; Meng et al., 2023; Zheng et al., 2023a). An emerging sub-family of distillation is the consistency-type models (Song et al., 2023; Song & Dhariwal, 2023; Kim et al., 2024), which exploit the fact that samples generated via different paths from the same initial noise should coincide.
@@
However, discrete diffusion models face a fundamental challenge when attempting to reduce the number of sampling steps. Conventional approaches use “product” models that treat each dimension independently as sampling distributions (Figure 1, center), since high-dimensional joint distributions are intractable. While this has been successful with hundreds of sampling steps, ignoring element-wise dependencies (which we refer to as dimensional correlations) causes non-negligible approximation errors (Figure 1). This ignoring is also pointed out in some concurrent works (Park et al., 2024; Liu et al., 2024; Xu et al., 2024).
@@
Experiments: Finally, we demonstrate that our approach is general and applicable to multiple settings. (1) On CIFAR-10 with a pixel-based discretized Gaussian diffusion, we substantially improve the sample quality metrics of the teacher model (Campbell et al., 2022) in few-step sampling. (2) On ImageNet class-conditional generation with masked diffusion, our method achieves a 2x speed-up while maintaining a comparable sample quality to the teacher model (Besnier & Chen, 2023). In addition, (3) on masked diffusion language modeling with OpenWebText, we show that Di4C can further distill a well-distilled model (Deschenaux & Gulcehre, 2024) by capturing dimensional correlations, without much harming of sampling diversity. These results consistently demonstrate that Di4C can effectively compress the sampling steps of discrete diffusion models while maintaining or improving generation quality.
@@ 3.2 Mixture models to capture dimensional correlations
$p_{s|t}^{\theta}(\bm{x}_{s}|\bm{x}_{t})=\mathbb{E}_{\lambda}\!\left[p_{s|t}^{\theta}(\bm{x}_{s}|\bm{x}_{t};\lambda)\right],\qquad\text{where}\ \ p_{s|t}^{\theta}(\bm{x}_{s}|\bm{x}_{t};\lambda)=\prod_{d=1}^{D}p_{s|t}^{\theta,d}(x_{s}^{d}|\bm{x}_{t};\lambda).$ (3)
@@
Indeed, we have $p(\bm{x})=\mathbb{E}_{\bm{z}\sim p}\!\left[\delta_{\bm{z}}(\bm{x})\right]$, where $\delta_{\bm{z}}$ is the delta distribution at $\bm{z}$, which is certainly a product distribution. Although the proof is not very informative, the assertion itself implies that the mixture model has sufficient expressive power to capture dimensional correlations. It should also be noted that sampling from this mixture model during the inference has almost no extra computational overhead compared with the conventional product model, since it just requires sampling and insertion of $\lambda$ (see Section F.1).
@@ 3.3 Consistency for distilling dimensional correlations
Figure 2: An illustration of how our loss functions work. Through $\mathcal{L}_{\mathrm{distil}}$ and $\mathcal{L}_{\mathrm{consis}}$, we distill the multiple teacher denoising steps into fewer steps of the student denoiser.
@@
$\mathcal{L}_{\mathrm{consis}}(\theta;\psi,r_{t},s,u,t):=\mathbb{E}_{\bm{x}_{t}\sim r_{t}}\!\left[D_{\mathrm{KL}}(p_{s|u}^{\theta}\circ p_{u|t}^{\psi}(\cdot|\bm{x}_{t})\,\|\,p_{s|t}^{\theta}(\cdot|\bm{x}_{t}))\right],$
@@ 4 Theoretical analysis
$d_{\mathrm{TV}}(p(\text{$1$-step student}),q_{0})$ $\leq d_{\mathrm{TV}}(p(\text{$1$-step student}),p(\text{$N$-step teacher}))+\mathcal{O}\!\left(1/N\right)$ (Theorem 1, upper bound)
@@ 4.1 Product models with multi-step sampling can approximate data distribution
Theorem 1 shows the capability and limitation of a dimensionally independent sampling scheme called analytical sampling (Sun et al., 2023) or Tweedie $\tau$-leaping (Lou et al., 2024; Ou et al., 2024), where we use a product-model denoiser $p_{s|t}(\bm{x}_{s}|\bm{x}_{t})=\prod_{d=1}^{D}p_{s|t}^{d}(x_{s}^{d}|\bm{x}_{t})$ approximating the true marginal as $p_{s|t}^{d}(x_{s}^{d}|\bm{x}_{t})\approx q_{s|t}^{d}(x_{s}^{d}|\bm{x}_{t})$. Although commonly used, there has been only empirical evidence for the overall efficiency of this dimensionally independent method. Note that Campbell et al. (2022) provides a guarantee for another dimensionally independent method called $\tau$-leaping.
@@ 4.2 Our losses can distill multi-step denoising models
$d_{\mathrm{TV}}\!\left(r_{0},\mathbb{E}_{\bm{x}_{T}\sim r_{T}}\!\left[p^{\theta}_{0|T}(\cdot|\bm{x}_{T})\right]\right)\leq\frac{1}{\sqrt{2}}\left(\mathcal{L}_{\mathrm{distil}}(\theta;\psi,r_{t_{1}},t_{1})^{1/2}+\sum_{n=1}^{N-1}\mathcal{L}_{\mathrm{consis}}(\theta;\psi,r_{t_{n+1}},0,t_{n},t_{n+1})^{1/2}\right).$ (6)
@@ 5 Experimental results
We evaluated our Di4C method in three different diffusion settings, each with distinct teacher models and state spaces. First, we examined continuous-time discrete-state diffusion with pixel-space representations using CIFAR-10, where we distilled from a well-trained U-Net teacher model (Section 5.1). Second, we explored masked diffusion on vector-quantized (VQ) space using ImageNet, working with a transformer-based teacher model designed for masked image generation (Section 5.2). Finally, we tested our approach on masked diffusion language models trained on OpenWebText, demonstrating Di4C’s effectiveness in distilling transformer-based diffusion language models (Section 5.3). These experiments, spanning different domains, architectures, and diffusion processes, showcase the broad applicability of our method while highlighting its consistent ability to achieve faster sampling.
@@ 5.1 Discretized Gaussian diffusion on pixel space
Table 1: Comparison of models on CIFAR-10 dataset. Fréchet inception distance (FID $\downarrow$) against training dataset and inception score (IS $\uparrow$) are shown in this order (FID / IS).
@@
Table 1 shows the results. The “hybrid” model used the student model for the first half of the denoising process and then switched to the teacher model for the remaining steps during inference. The student substantially improved the metrics compared with the teacher in 10-step sampling, while the benefits of our method diminished as the number of steps grew. In contrast, the hybrid model was the best at 20 steps and on par with the teacher in 40-step FID. We hypothesize that Di4C is particularly effective when using fewer sampling steps, as this is where capturing complex dimensional correlations is crucial (Theorem 1).
@@ 5.2 Masked image modeling on VQ space
Figure 3: FID-IS curves of the $4/8$-step teacher and $4$-step Di4C models on ImageNet $256\times 256$ when varying CFG coefficients. The arrows connect the experimental results (dots) at each CFG coefficient in ascending order.
Figure 3 shows FID-IS curves of 4-step sampling with various CFG coefficients: $w_{\mathrm{cfg}}\in\{5,6,7,8,9,10,11,12\}$ for 4 steps and $w_{\mathrm{cfg}}\in\{2,2.5,3,3.5,4\}$ for 8 steps (teacher only). In the figure, di4c represents a model trained using the standard Di4C method, while di4c-d represents a model trained with an additional datapoint loss (see Section F.3.4 for details). As shown in the figure, the FID-IS performance of the 4-step Di4C models closely matched that of the 8-step sampling of the teacher model. This result indicates that Di4C can achieve an approximate 2x speed-up in the sampling process. See Section F.3 for further details.
5.3 Masked diffusion language models
Finally, we examined Di4C in language modeling. As teacher models, we adopted two versions of pretrained SDTT models (Deschenaux & Gulcehre, 2024, with a KL target), which we refer to as sdtt-6 and sdtt-7. They were obtained after respectively 6 and 7 rounds of distillation of a masked diffusion language model (MDLM; Sahoo et al., 2024) trained on the OpenWebText dataset (Gokaslan & Cohen, 2019). As a forward process, they use a masked diffusion with a GPT-2 tokenizer (Radford et al., 2019), which is essentially the same as the one explained in Section 5.2, while we had $D=1024$ and $\lvert\mathcal{S}^{*}\rvert=50257$ this time. In the mixture modeling for Di4C, we added one token from $\lambda\sim\mathrm{Unif}([0,1])$ to the transformer, similarly to the previous section. See Section F.4 for details.
@@ 5.3 Masked diffusion language models
Figure 4: Comparison of SDTT checkpoints (Deschenaux & Gulcehre, 2024) and their Di4C distillations.
@@
Figure 4(a) shows a performance comparison for unconditional generation. Each model generated 1024 samples of 1024 tokens, with the number of sampling steps in $\{8,16,32,64,128\}$. Following the previous work (Lou et al., 2024; Sahoo et al., 2024; Deschenaux & Gulcehre, 2024), we used the GPT-2-large model (Radford et al., 2019) to evaluate the generative perplexity. The results show that applying one round of Di4C lead to a similar level of improvement as applying one round of SDTT (compare sdtt-6 + di4c with sdtt-7). Also, two rounds of Di4C (+ di4c2) showed speed-ups of more than 2x to match the teacher’s performance of 64 or 128 steps.
@@ 6 Related work
For faster sample generation in discrete diffusion models, Park et al. (2024) propose a post-hoc optimization of sampling schedules and outperforms the uniform partitioning of $[0,T]$. To speed up MDLMs (Sahoo et al., 2024), SDTT (Deschenaux & Gulcehre, 2024) gathers the logits at unmasked tokens from a pretrained teacher MDLM throughout the denoising process, achieving a speed-up of approximately 32x compared with the teacher model. In addition, the confidence-based sampling by MaskGIT (Chang et al., 2022) is essentially a heuristic to enable faster sampling in masked diffusions. These approaches can be combined with Di4C; we indeed have demonstrated it for the latter two.

### v4

@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b) have demonstrated excellent performance in generative modeling, particularly for continuous data such as images (Nichol et al., 2021; Rombach et al., 2022; Saharia et al., 2022), audio (Kong et al., 2021; Chen et al., 2021; Evans et al., 2024), and video (Harvey et al., 2022; Ho et al., 2022; Blattmann et al., 2023). Recent advancements in diffusion models often outperform traditional generative models, such as variational autoencoders (VAEs, Kingma & Welling, 2014; Higgins et al., 2017; Zhao et al., 2019) and generative adversarial networks (GANs, Goodfellow et al., 2014), in terms of sample quality and the controllability of the generated results. Furthermore, diffusion models are not limited to learning continuous data; they can also be applied to discrete or categorical data with modifications (Hoogeboom et al., 2021; Austin et al., 2021) and offer a promising approach for discrete generative modeling (Gu et al., 2022; Lou et al., 2024). Such discrete diffusion models are the main topic of this paper.
A notable drawback of diffusion models is their slow sampling speed due to requiring many sampling steps (Xiao et al., 2022; Zhang & Chen, 2023). In continuous domains, various approaches have been proposed to reduce the number of steps, including well-designed forward processes (Song et al., 2021a) and fast solvers of stochastic/ordinary differential equations (SDEs/ODEs, Lu et al., 2022a, b; Zheng et al., 2023b). Another notable approach is knowledge distillation, which significantly reduces the number of sampling steps compared with earlier attempts by compressing pretrained diffusion models into single- or few-step generative models (Luhman & Luhman, 2021; Salimans & Ho, 2022; Meng et al., 2023; Zheng et al., 2023a). An emerging sub-family of distillation is the consistency-type models (Song et al., 2023; Song & Dhariwal, 2023; Kim et al., 2024), which exploit the fact that samples generated via different paths from the same initial noise should coincide.
@@
However, discrete diffusion models face a fundamental challenge when attempting to reduce the number of sampling steps. Conventional approaches use “product” models that treat each dimension independently as sampling distributions (Figure 1, center), since high-dimensional joint distributions are intractable. While this has been successful with hundreds of sampling steps, ignoring element-wise dependencies (which we refer to as dimensional correlations) causes non-negligible approximation errors (Figure 1). This ignoring is also pointed out in some concurrent works (Park et al., 2025; Liu et al., 2024; Xu et al., 2025).
@@
Experiments: Finally, we demonstrate that our approach is general and applicable to multiple settings. (1) On CIFAR-10 with a pixel-based discretized Gaussian diffusion, we substantially improve the sample quality metrics of the teacher model (Campbell et al., 2022) in few-step sampling. (2) On ImageNet class-conditional generation with masked diffusion, our method achieves a 2x speed-up while maintaining a comparable sample quality to the teacher model (Besnier & Chen, 2023). In addition, (3) on masked diffusion language modeling with OpenWebText, we show that Di4C can further distill a well-distilled model (Deschenaux & Gulcehre, 2025) by capturing dimensional correlations, without much harming of sampling diversity. These results consistently demonstrate that Di4C can effectively compress the sampling steps of discrete diffusion models while maintaining or improving generation quality.
@@ 3.2 Mixture models to capture dimensional correlations
$p_{s|t}^{\theta}(\bm{x}_{s}|\bm{x}_{t})=\mathbb{E}_{\lambda}\!\left[p_{s|t}^{\theta}(\bm{x}_{s}|\bm{x}_{t};\lambda)\right],$ (3)
$\text{where}\quad p_{s|t}^{\theta}(\bm{x}_{s}|\bm{x}_{t};\lambda)=\prod_{d=1}^{D}p_{s|t}^{\theta,d}(x_{s}^{d}|\bm{x}_{t};\lambda).$
@@
Indeed, we have $p(\bm{x})=\mathbb{E}_{\bm{z}\sim p}\!\left[\delta_{\bm{z}}(\bm{x})\right]$, where $\delta_{\bm{z}}$ is the delta distribution at $\bm{z}$, which is certainly a product distribution. Although the proof is not very informative, the assertion itself implies that the mixture model has sufficient expressive power to capture dimensional correlations. It should also be noted that sampling from this mixture model during the inference has almost no extra computational overhead compared with the conventional product model, since it just requires sampling and insertion of $\lambda$ (see Section 5.4).
@@ 3.3 Consistency for distilling dimensional correlations
$\mathcal{L}_{\mathrm{consis}}(\theta;\psi,r_{t},s,u,t)$
$:=\mathbb{E}_{\bm{x}_{t}\sim r_{t}}\!\left[D_{\mathrm{KL}}(p_{s|u}^{\theta}\circ p_{u|t}^{\psi}(\cdot|\bm{x}_{t})\,\|\,p_{s|t}^{\theta}(\cdot|\bm{x}_{t}))\right],$
@@
Figure 2: Illustration of how our loss functions work. Through $\mathcal{L}_{\mathrm{distil}}$ and $\mathcal{L}_{\mathrm{consis}}$, we distill multiple teacher denoising steps into fewer steps of student denoiser.
@@ 4 Theoretical analysis
[…]
$\leq d_{\mathrm{TV}}(p(\text{$1$-step student}),p(\text{$N$-step teacher}))+\mathcal{O}\!\left(1/N\right)$
@@ 4.1 Product models with multi-step sampling can approximate data distribution
Theorem 1 shows the capability and limitation of a dimensionally independent sampling scheme called analytical sampling (Sun et al., 2023) or Tweedie $\tau$-leaping (Lou et al., 2024; Ou et al., 2025), where we use a product-model denoiser $p_{s|t}(\bm{x}_{s}|\bm{x}_{t})=\prod_{d=1}^{D}p_{s|t}^{d}(x_{s}^{d}|\bm{x}_{t})$ approximating the true marginal as $p_{s|t}^{d}(x_{s}^{d}|\bm{x}_{t})\approx q_{s|t}^{d}(x_{s}^{d}|\bm{x}_{t})$. Although commonly used, there has been only empirical evidence for the overall efficiency of this dimensionally independent method. Note that Campbell et al. (2022) provides a guarantee for another dimensionally independent method called $\tau$-leaping.
@@ 4.2 Our losses can distill multi-step denoising models
$d_{\mathrm{TV}}\!\left(r_{0},\mathbb{E}_{\bm{x}_{T}\sim r_{T}}\!\left[p^{\theta}_{0|T}(\cdot|\bm{x}_{T})\right]\right)$
$\leq\frac{1}{\sqrt{2}}\Bigl{(}\mathcal{L}_{\mathrm{distil}}(\theta;\psi,r_{t_{1}},t_{1})^{1/2}$
$\quad\qquad+\sum_{n=1}^{N-1}\mathcal{L}_{\mathrm{consis}}(\theta;\psi,r_{t_{n+1}},0,t_{n},t_{n+1})^{1/2}\Bigr{)}.$ (6)
@@ 5 Experimental results
We evaluated our Di4C method in three different diffusion settings, each with distinct teacher models and state spaces. First, we examined continuous-time discrete-state diffusion with pixel-space representations using CIFAR-10, where we distilled from a well-trained U-Net teacher model (Section 5.1). Second, we explored masked diffusion on vector-quantized (VQ) space using ImageNet, working with a transformer-based teacher model designed for masked image generation (Section 5.2). Third, we tested our approach on masked diffusion language models trained on OpenWebText, demonstrating Di4C’s effectiveness in distilling transformer-based diffusion language models (Section 5.3). Finally, we demonstrate that our mixture modeling causes minimal latency overhead in all the above experiments (Section 5.4). These experimental results, spanning different domains, architectures, and diffusion processes, showcase the broad applicability of our method while highlighting its consistent ability to achieve faster sampling.
@@ 5.1 Discretized Gaussian diffusion on pixel space
Table 1: Comparison of models on CIFAR-10 dataset. Fréchet inception distance (FID $\downarrow$) against training dataset and inception score (IS $\uparrow$) measured with 50,000 samples are shown in this order (FID / IS).
@@
Table 1 shows the results. The “hybrid” model used the student model for the first half of the denoising process and then switched to the teacher model for the remaining steps during inference. The student substantially improved the metrics compared with the teacher in 10-step sampling, while the benefits of our method diminished as the number of steps grew. In contrast, the hybrid model was the best at 20 steps and on par with the teacher in 40-step FID. We hypothesize that Di4C is particularly effective when using fewer sampling steps, as this is where capturing complex dimensional correlations is crucial (Theorem 1). We also give additional results for different sampling steps in Table 4.
Table 2: Comparison of teacher and student with PC steps. FID against CIFAR-10 training dataset measured with 10,000 generated samples are displayed (so there is discrepancy from Table 1). Number of function evaluations (NFEs) is adjusted to 20, with varying budget allocations between denoising and PC steps.
NFE 6+7+7 10+5+5 14+3+3 20+0+0
teacher 57.57 32.74 21.32 14.42
student 44.57 22.63 14.25 11.81
In Table 2, we also compare the teacher and student models in combination with the predictor-corrector (PC, Campbell et al., 2022) steps. The notation $m$+$n$+$n$ indicates that we used $n$ corrector steps before each of the final two out of $m$ denoising steps, imitating the approach in Campbell et al. (2022). For each setting, the best PC step-size among $\{10^{-3},10^{-4},10^{-5}\}$ was chosen. Since a PC step requires an additional network evaluation (as expensive as one step of denoising), we adjusted the total NFE in the table. While adding PC steps shows improvement upon the vanilla $m$-step denoising (see also Table 4) in both teacher and student models, allocating all the NFE budget to denoising steps performs better in this small NFE regime; this observation aligns with Campbell et al. (2022, Fig. 4).
@@ 5.2 Masked image modeling on VQ space
Figure 3: FID-IS curves of $4/8$-step teacher and $4$-step Di4C models on ImageNet $256\times 256$ when varying CFG coefficients. Arrows connect experimental results (dots) at each CFG coefficient in ascending order.
@@
Figure 4: Comparison of SDTT checkpoints (Deschenaux & Gulcehre, 2025) and their Di4C distillations.
Figure 3 shows FID-IS curves of 4-step sampling with various CFG coefficients: $w_{\mathrm{cfg}}\in\{5,6,7,8,9,10,11,12\}$ for 4 steps and $w_{\mathrm{cfg}}\in\{2,2.5,3,3.5,4\}$ for 8 steps (teacher only). In the figure, di4c represents a model trained using the standard Di4C method, while di4c-d represents a model trained with an additional datapoint loss (see Section F.2.4 for details). As shown in the figure, the FID-IS performance of the 4-step Di4C models closely matched that of the 8-step sampling of the teacher model. This result indicates that Di4C can achieve an approximate 2x speed-up in the sampling process.
See Section F.2 for further details of this experiment. In particular, Table 8 shows the detailed quantitative performance (including precision and recall) of each model at its best CFG coefficient.
5.3 Masked diffusion language models
Finally, we examined Di4C in language modeling. As teacher models, we adopted two versions of pretrained SDTT models (Deschenaux & Gulcehre, 2025, with a KL target), which we refer to as sdtt-6 and sdtt-7. They were obtained after respectively 6 and 7 rounds of distillation of a masked diffusion language model (MDLM, Sahoo et al., 2024. See also its concurrent works: Shi et al. 2024; Ou et al. 2025) trained on the OpenWebText dataset (Gokaslan & Cohen, 2019). As a forward process, they use a masked diffusion with a GPT-2 tokenizer (Radford et al., 2019), which is essentially the same as the one explained in Section 5.2, while we had $D=1024$ and $\lvert\mathcal{S}^{*}\rvert=50257$ this time. In the mixture modeling for Di4C, we added one token from $\lambda\sim\mathrm{Unif}([0,1])$ to the transformer, similarly to the previous section. See Section F.3 for details.
@@ 5.3 Masked diffusion language models
Figure 4(a) shows a performance comparison for unconditional generation. Each model generated 1024 samples of 1024 tokens, with the number of sampling steps in $\{8,16,32,64,128\}$. Following the previous work (Lou et al., 2024; Sahoo et al., 2024; Deschenaux & Gulcehre, 2025), we used the GPT-2-large model (Radford et al., 2019) to evaluate the generative perplexity. The results show that applying one round of Di4C lead to a similar level of improvement as applying one round of SDTT (compare sdtt-6 + di4c with sdtt-7). Also, two rounds of Di4C (+ di4c2) showed speed-ups of more than 2x to match the teacher’s performance of 64 or 128 steps.
@@
5.4 Latency overhead of mixture modeling
Experiment # steps teacher (sec.) student (sec.)
CIFAR-10 10 0.5515$\pm$0.0024 0.5786$\pm$0.0017
ImageNet 4 2.0741$\pm$0.0035 2.0734$\pm$0.0043
OpenWebText 16 3.3409$\pm$0.0417 3.4817$\pm$0.0730
@@ 6 Related work
For faster sample generation in discrete diffusion models, Park et al. (2025) propose a post-hoc optimization of sampling schedules and outperforms the uniform partitioning of $[0,T]$. To speed up MDLMs (Sahoo et al., 2024); SDTT (Deschenaux & Gulcehre, 2025) gathers the logits at unmasked tokens from a pretrained teacher MDLM throughout the denoising process, achieving a speed-up of approximately 32x compared with the teacher model. In addition, the confidence-based sampling by MaskGIT (Chang et al., 2022) is essentially a heuristic to enable faster sampling in masked diffusions. These approaches can be combined with Di4C; we indeed have demonstrated it for the latter two.